# DECEPTICON: How Dark Patterns Manipulate Web Agents

### v1

@@ 3 The Decepticon Environment
Task structure. A single task comprises three components: (1) a verifiable objective, e.g., “buy a bouquet of flowers under \$30”; (2) a desired goal state, e.g., the order confirmation page for the cheapest qualifying bouquet; and (3) a dark pattern, e.g., a pre-checked premium shipping option. A task is considered complete when the agent terminates its trajectory, either by sending a task completion signal (§C) or by reaching the maximum step count of 15.
@@
We construct 600 synthetic dark pattern tasks using an adversarial generation pipeline. First, we generate base website UIs in a single pass for several common web navigation tasks (e.g., e-commerce, event/subscription booking, information retrieval) using Gemini-2.5-Flash (comanici2025gemini25pushingfrontier). Next, dark patterns are generated based on visual and textual descriptions from Mathur_2019 and Nouwens_2020 using Gemini-2.5-Pro and an agentic scaffold. An agent then naively attempts the task in the generated environment to verify whether the task is solvable; its results serve as reward signals to increase the difficulty of the dark pattern implementation in the next iteration. For our experiments, we run only a single iteration of this generate-and-test loop to ensure the dark patterns are not overfitted to the agent’s behavior. Finally, human verification ensures that the dark pattern is correctly implemented, the task is solvable, and is not redundant with existing patterns—almost 70% of generated dark pattern tasks were filtered this way. The full pipeline is described in Appendix §D. Because dark pattern examples are sourced directly from prior works, this approach ensures balanced representation across dark pattern categories and classes, enabling controlled experimentation with standardized implementations. Class distribution is reported in Appendix §D. Each generated task is designed to isolate specific dark pattern mechanisms while maintaining realistic web interface.
@@
We collect 100 real-world dark pattern instances through an agent-driven web scraping approach, which we discuss in greater detail in Appendix §D. Starting with a curated set of live web pages drawn from (Mathur_2019), the Ahrefs database of popular websites (Linehan2025EcommerceStartups), and prior collections of documented dark pattern instances (Nouwens_2020; Luguri2019Shining), we deploy agents across these sites using seed objectives based on common user workflows. At each trajectory step, an LLM-based detector agent identifies potential dark patterns, followed by human validation to confirm authenticity and relevance, and subsequent website archiving. This methodology captures the diversity and complexity of dark patterns as they appear in real-world production web environments.
@@ 4.4 Human Benchmark
Under these conditions, human participants achieve an overall SR of 81.0% on the generated split and 80.8% on the in-the-wild split, with corresponding DP rates of 31.0% and 33.4%, respectively (see Table 2(a)). This indicates that while humans are susceptible to dark patterns, they are significantly more adept at avoiding them compared to LLM-driven agents, which exhibit DP rates exceeding 70% across both splits. The gap in performance highlights the deficiency of current web agents in handling deceptive web design, underscoring the need for improved robustness mechanisms. Further details on the configuration of the human study, as well as a taxonomy-category breakdown of results, are presented in Appendix J.
@@ 6 Defenses Against Dark Patterns
We test the four (Simple) agents (GPT-4o, GPT-5, Gemini-2.5-Flash, Gemini-2.5-Pro) (openai2024gpt4ocard; openai2025gpt5; comanici2025gemini25pushingfrontier) with both defense mechanisms, retaining the same environment, tasks, and configuration from RQ1 $(maxsteps=15,temperature=0)$, across the entire generated task suite of $(N=600)$ tasks, and compare their performance to the baseline agents from RQ1. Further details on the implementation of these defenses are described in Appendix § C.2.
@@
CoT of defense failures reveals challenges in both detection and reasoning. Inspection of agent CoTs show that errors of the third category (erroneous reasoning, as described in Section 5) predominate, indicating that reasoning misalignment is core to the weakness of LLMs to dark patterns. For ICP, the agent ignores the effects of the dark pattern: “[…] the terms and conditions. Of course, nobody reads these, but I’ll humor them. Ah, here’s the magic button: "TLDR: ACCEPT" […] zeroing out any need for, you know, actually *reading* anything”, or acknowledges the dark pattern, but chooses to proceed with it regardless, with the CoT indicating that it is necessary or helpful (we list this example in Appendix § K). For guardrail model agents, the higher detection rate of dark patterns does not preclude erroneous reasoning, with CoTs indicating ignorance of the dark pattern or confusion of instruction authority; these are also presented in Appendix § K.
@@ Appendix A Open-Source Agent Results
Appendix B Different URL Ablation
@@
Appendix C Additional Agent Details
C.1 Simple Agent System
The Simple agent system prompt is adapted from the WebVoyager agent (he2024webvoyager), with modifications to better suit the dark pattern evaluation tasks. Namely, options to perform a Google search and download files were removed, as these actions were not relevant to the evaluation scenarios. The full prompt is provided below.
@@ C.1 Simple Agent System
C.2 In-Context Prompting Defense System Prompt
@@ 6. Forced Action
C.3 Guardrail Model Defense Prompt
The following defense prompts are used as the guardrail model defense prompt. As implemented, the agent is provided with the base Simple agent prompt (Appendix §C.1) along with one of the following defense prompts, depending on the category of dark pattern detected by the detector LLM; during testing, the detector was assumed to be perfect, with the appropriate prompt always provided. This is not necessarily reflective of real-world performance, but allows us to isolate the effectiveness of the defense prompts themselves. Each prompt describes the relevant dark pattern category, provides examples of how it may manifest, and offers specific instructions on how to avoid being influenced by the pattern.
@@ 1. Forced Enrollment: Requiring users to sign up or agree to
Appendix D Dataset Compilation Details
D.1 Notes on Statistics
@@
D.2 Data Collection Pipeline Overview
@@
D.3 Generated Website Adversarial Generation Pipeline
@@
D.4 Generated Website Examples
@@
D.5 Task Distribution by Category - Generated
@@
D.6 In-the-Wild Website Examples
Collection of in-the-wild tasks was possible via several alternative collection methods, but the authors found that a crawler-driven collection was the most efficient and scalable. Manual collection was possible, but required significant human effort to identify and verify dark patterns, while crowdsourced collection was attempted but proved difficult to verify and required significant quality control. Therefore, we collected a list of seed websites from lists of known dark pattern-containing websites (Mathur_2019), and then crawled these websites using a breadth-first search strategy, following links up to a depth of 3 from the seed URLs. We then used a Gemini-2.5-Flash based classifier to identify pages likely to contain dark patterns  [^3] The prompt used in in-context defense was modified and used as the system prompt for the classifier, filtered out duplicates and non-HTML content, and manually verified the resultant pages to ensure the presence of dark patterns. Finally, we induced tasks on these pages by running a Gemini-2.5-Flash agent to modify tasks (taken from a list of seed tasks (he2024webvoyager)) with content from the website; then manually verified these tasks to ensure they were solvable and contained dark patterns. We archived the sites using wget to ensure reproducibility, serving the interactable archives during all in-the-wild tests. This process yielded 100 verified in-the-wild tasks.
@@ D.6 In-the-Wild Website Examples
D.7 Task Distribution by Category - In-the-Wild
@@
Appendix E Additional RQ2 Results
@@
E.1 Model Size Scaling Results
Table 6 shows the complete results for the Qwen2.5-VL (bai2025qwen25vltechnicalreport) model family across different parameter sizes, demonstrating how larger models achieve higher task success rates but exhibit increased vulnerability to dark patterns.
@@ E.1 Model Size Scaling Results
E.2 Reasoning Tokens Scaling Results
Table 7 presents the results for Gemini-2.5-Flash (comanici2025gemini25pushingfrontier) across different thinking budget allocations.
@@ E.2 Reasoning Tokens Scaling Results
E.3 Reasoning Effort Scaling Results
Table 8 shows the OpenAI o3 (openai2025o3systemcard) model performance across different reasoning effort levels.
@@ E.3 Reasoning Effort Scaling Results
Appendix F Additional RQ3 Results
@@
F.1 Defense Mechanism Effectiveness
Table 9 presents the comprehensive results for all tested models (GPT-4o (openai2024gpt4ocard), GPT-5 (openai2025gpt5), Gemini-2.5-Flash, Gemini-2.5-Pro (comanici2025gemini25pushingfrontier)) across the three conditions: baseline (no defense), In-Context Prompting, and guardrail models.
@@ F.1 Defense Mechanism Effectiveness
Appendix G Independent Guardrail Model Details
We perform guardrail model verification as described in RQ3 by using a secondary LLM to identify dark patterns present in the environment. We perform an independent study of the effectiveness of the LLM in detecting the dark patterns present in the environment, to qualify how well LLMs can identify dark patterns from screenshots and text. We test the four model configurations used in RQ1 within the Simple agent configuration (GPT-4o, GPT-5, Gemini-2.5-Flash, Gemini-2.5-Pro) (openai2024gpt4ocard; comanici2025gemini25pushingfrontier), and prompt them with the same information provided to the agent in RQ1: a screenshot of the current environment state, the text content of the page, and any other observations provided by the agent scaffolding (e.g., current URL). The LLM is prompted to identify whether a dark pattern is present in the environment, and if so, which element corresponds to the dark pattern; if no dark pattern is present, the LLM is instructed to respond with "No dark pattern detected". The full prompt used is provided in Appendix § C.2.
@@ Appendix G Independent Guardrail Model Details
Appendix H Comparison with Related Benchmarks and Attack Settings
Table 11 provides a comparison of Decepticon with other web agent benchmarks (WebArena (zhou2023webarena), VisualWebArena (koh2024visualwebarena), WorkArena (drouin2024workarenacapablewebagents), BrowseComp (wei2025browsecompsimplechallengingbenchmark)) and agent-specific attack environments (pop-up attacks (zhang2025attackingvisionlanguagecomputeragents), environmental injection attacks (liao2025eiaenvironmentalinjectionattack)) across key dimensions. Notably, Decepticon is the only benchmark that focuses on human-targeted adversarial elements, provides non-adversarial control pages for direct comparison, ensures archival determinism for reproducible offline evaluation, enforces avoidability constraints on dark patterns, includes a structured category taxonomy for adversarial patterns, employs dual metrics for success rate and dark pattern effectiveness, and systematically evaluates defense mechanisms. This comprehensive approach enables a more thorough understanding of dark pattern impacts on web agents compared to existing benchmarks and attack settings.
@@ Appendix H Comparison with Related Benchmarks and Attack Settings
Appendix I Modality Ablation
This section presents an ablation study on the input modalities, in order to quantify the contribution of different input modalities to dark pattern effectiveness. We conduct experiments using GLM-4.5V (vteam2025glm45vglm41vthinkingversatilemultimodal) on the Generated split of Decepticon, varying the input modalities provided to the model. We consider three configurations: (1) Accessibility Tree (a11y): only the accessibility tree representation of the webpage is provided as input; (2) Coordinates (coordinates): only the coordinates of clickable elements on the page are provided as input; (3) Screenshot + JS-based Set of Marks (SoM): both the screenshot of the webpage and the JS-based Set of Marks representation are provided as input. The results are summarized in Table 13.
@@ Appendix I Modality Ablation
Appendix J Details on Human Study of Decepticon Tasks
@@
Appendix K Examples of Failures in Defenses
@@
K.1 In-Context Prompting Failure Examples
@@
K.2 Guardrail Failure Examples

### v2

@@ 3 The Decepticon Environment
Task structure. A single task comprises three components: (1) a verifiable objective, e.g., “buy a bouquet of flowers under \$30”; (2) a desired goal state, e.g., the order confirmation page for the cheapest qualifying bouquet; and (3) a dark pattern, e.g., a pre-checked premium shipping option. A task is considered complete when the agent terminates its trajectory, either by sending a task completion signal (§D) or by reaching the maximum step count of 15.
@@
We construct 600 synthetic dark pattern tasks using an adversarial generation pipeline. First, we generate base website UIs in a single pass for several common web navigation tasks (e.g., e-commerce, event/subscription booking, information retrieval) using Gemini-2.5-Flash (comanici2025gemini25pushingfrontier). Next, dark patterns are generated based on visual and textual descriptions from Mathur_2019 and Nouwens_2020 using Gemini-2.5-Pro and an agentic scaffold. An agent then naively attempts the task in the generated environment to verify whether the task is solvable; its results serve as reward signals to increase the difficulty of the dark pattern implementation in the next iteration. For our experiments, we run only a single iteration of this generate-and-test loop to ensure the dark patterns are not overfitted to the agent’s behavior. Finally, human verification ensures that the dark pattern is correctly implemented, the task is solvable, and is not redundant with existing patterns—almost 70% of generated dark pattern tasks were filtered this way. The full pipeline is described in Appendix §E. Because dark pattern examples are sourced directly from prior works, this approach ensures balanced representation across dark pattern categories and classes, enabling controlled experimentation with standardized implementations. Class distribution is reported in Appendix §E. Each generated task is designed to isolate specific dark pattern mechanisms while maintaining realistic web interface.
@@
We collect 100 real-world dark pattern instances through an agent-driven web scraping approach, which we discuss in greater detail in Appendix §E. Starting with a curated set of live web pages drawn from (Mathur_2019), the Ahrefs database of popular websites (Linehan2025EcommerceStartups), and prior collections of documented dark pattern instances (Nouwens_2020; Luguri2019Shining), we deploy agents across these sites using seed objectives based on common user workflows. At each trajectory step, an LLM-based detector agent identifies potential dark patterns, followed by human validation to confirm authenticity and relevance, and subsequent website archiving. This methodology captures the diversity and complexity of dark patterns as they appear in real-world production web environments.
@@ 4.4 Human Benchmark
Under these conditions, human participants achieve an overall SR of 81.0% on the generated split and 80.8% on the in-the-wild split, with corresponding DP rates of 31.0% and 33.4%, respectively (see Table 2(a)). This indicates that while humans are susceptible to dark patterns, they are significantly more adept at avoiding them compared to LLM-driven agents, which exhibit DP rates exceeding 70% across both splits. The gap in performance highlights the deficiency of current web agents in handling deceptive web design, underscoring the need for improved robustness mechanisms. Further details on the configuration of the human study, as well as a taxonomy-category breakdown of results, are presented in Appendix K.
@@ 6 Defenses Against Dark Patterns
We test the four (Simple) agents (GPT-4o, GPT-5, Gemini-2.5-Flash, Gemini-2.5-Pro) (openai2024gpt4ocard; openai2025gpt5; comanici2025gemini25pushingfrontier) with both defense mechanisms, retaining the same environment, tasks, and configuration from RQ1 $(maxsteps=15,temperature=0)$, across the entire generated task suite of $(N=600)$ tasks, and compare their performance to the baseline agents from RQ1. Further details on the implementation of these defenses are described in Appendix § D.2.
@@
CoT of defense failures reveals challenges in both detection and reasoning. Inspection of agent CoTs show that errors of the third category (erroneous reasoning, as described in Section 5) predominate, indicating that reasoning misalignment is core to the weakness of LLMs to dark patterns. For ICP, the agent ignores the effects of the dark pattern: “[…] the terms and conditions. Of course, nobody reads these, but I’ll humor them. Ah, here’s the magic button: "TLDR: ACCEPT" […] zeroing out any need for, you know, actually *reading* anything”, or acknowledges the dark pattern, but chooses to proceed with it regardless, with the CoT indicating that it is necessary or helpful (we list this example in Appendix § L). For guardrail model agents, the higher detection rate of dark patterns does not preclude erroneous reasoning, with CoTs indicating ignorance of the dark pattern or confusion of instruction authority; these are also presented in Appendix § L.
@@ Appendix A Open-Source Agent Results
Appendix B Closed-Source Agent Results
This section presents evaluation results for closed-source CUA agents on the Decepticon benchmark. We evaluate the performance and dark pattern susceptibility of closed-source CUA agents. The results are summarized in Table 4.
Table 4: Performance of closed-source CUA agents on the Decepticon benchmark.
[…]
claude-sonnet-4-20250514 (anthropic_cua) 60 34.5 46.6 27.6
computer-use-preview (oai_cua) 60 23.3 41.7 38.3
Appendix C Different URL Ablation
@@
Appendix D Additional Agent Details
D.1 Simple Agent System
The Simple agent system prompt is adapted from the WebVoyager agent [he2024webvoyager], with modifications to better suit the dark pattern evaluation tasks. Namely, options to perform a Google search and download files were removed, as these actions were not relevant to the evaluation scenarios. The full prompt is provided below.
@@ D.1 Simple Agent System
D.2 In-Context Prompting Defense System Prompt
@@ 6. Forced Action
D.3 Guardrail Model Defense Prompt
The following defense prompts are used as the guardrail model defense prompt. As implemented, the agent is provided with the base Simple agent prompt (Appendix §D.1) along with one of the following defense prompts, depending on the category of dark pattern detected by the detector LLM; during testing, the detector was assumed to be perfect, with the appropriate prompt always provided. This is not necessarily reflective of real-world performance, but allows us to isolate the effectiveness of the defense prompts themselves. Each prompt describes the relevant dark pattern category, provides examples of how it may manifest, and offers specific instructions on how to avoid being influenced by the pattern.
@@ 1. Forced Enrollment: Requiring users to sign up or agree to
Appendix E Dataset Compilation Details
E.1 Notes on Statistics
@@
E.2 Data Collection Pipeline Overview
@@
E.3 Generated Website Adversarial Generation Pipeline
@@
E.4 Generated Website Examples
@@
E.5 Task Distribution by Category - Generated
@@
E.6 In-the-Wild Website Examples
Collection of in-the-wild tasks was possible via several alternative collection methods, but the authors found that a crawler-driven collection was the most efficient and scalable. Manual collection was possible, but required significant human effort to identify and verify dark patterns, while crowdsourced collection was attempted but proved difficult to verify and required significant quality control. Therefore, we collected a list of seed websites from lists of known dark pattern-containing websites [Mathur_2019], and then crawled these websites using a breadth-first search strategy, following links up to a depth of 3 from the seed URLs. We then used a Gemini-2.5-Flash based classifier to identify pages likely to contain dark patterns  [^3] The prompt used in in-context defense was modified and used as the system prompt for the classifier, filtered out duplicates and non-HTML content, and manually verified the resultant pages to ensure the presence of dark patterns. Finally, we induced tasks on these pages by running a Gemini-2.5-Flash agent to modify tasks (taken from a list of seed tasks [he2024webvoyager]) with content from the website; then manually verified these tasks to ensure they were solvable and contained dark patterns. We archived the sites using wget to ensure reproducibility, serving the interactable archives during all in-the-wild tests. This process yielded 100 verified in-the-wild tasks.
@@ E.6 In-the-Wild Website Examples
E.7 Task Distribution by Category - In-the-Wild
@@
Appendix F Additional RQ2 Results
@@
F.1 Model Size Scaling Results
Table 7 shows the complete results for the Qwen2.5-VL [bai2025qwen25vltechnicalreport] model family across different parameter sizes, demonstrating how larger models achieve higher task success rates but exhibit increased vulnerability to dark patterns.
@@ F.1 Model Size Scaling Results
F.2 Reasoning Tokens Scaling Results
Table 8 presents the results for Gemini-2.5-Flash [comanici2025gemini25pushingfrontier] across different thinking budget allocations.
@@ F.2 Reasoning Tokens Scaling Results
F.3 Reasoning Effort Scaling Results
Table 9 shows the OpenAI o3 [openai2025o3systemcard] model performance across different reasoning effort levels.
@@ F.3 Reasoning Effort Scaling Results
Appendix G Additional RQ3 Results
@@
G.1 Defense Mechanism Effectiveness
Table 10 presents the comprehensive results for all tested models (GPT-4o [openai2024gpt4ocard], GPT-5 [openai2025gpt5], Gemini-2.5-Flash, Gemini-2.5-Pro [comanici2025gemini25pushingfrontier]) across the three conditions: baseline (no defense), In-Context Prompting, and guardrail models.
@@ G.1 Defense Mechanism Effectiveness
Appendix H Independent Guardrail Model Details
We perform guardrail model verification as described in RQ3 by using a secondary LLM to identify dark patterns present in the environment. We perform an independent study of the effectiveness of the LLM in detecting the dark patterns present in the environment, to qualify how well LLMs can identify dark patterns from screenshots and text. We test the four model configurations used in RQ1 within the Simple agent configuration (GPT-4o, GPT-5, Gemini-2.5-Flash, Gemini-2.5-Pro) [openai2024gpt4ocard, comanici2025gemini25pushingfrontier], and prompt them with the same information provided to the agent in RQ1: a screenshot of the current environment state, the text content of the page, and any other observations provided by the agent scaffolding (e.g., current URL). The LLM is prompted to identify whether a dark pattern is present in the environment, and if so, which element corresponds to the dark pattern; if no dark pattern is present, the LLM is instructed to respond with "No dark pattern detected". The full prompt used is provided in Appendix § D.2.
@@ Appendix H Independent Guardrail Model Details
Appendix I Comparison with Related Benchmarks and Attack Settings
Table 12 provides a comparison of Decepticon with other web agent benchmarks (WebArena [zhou2023webarena], VisualWebArena [koh2024visualwebarena], WorkArena [drouin2024workarenacapablewebagents], BrowseComp [wei2025browsecompsimplechallengingbenchmark]) and agent-specific attack environments (pop-up attacks [zhang2025attackingvisionlanguagecomputeragents], environmental injection attacks [liao2025eiaenvironmentalinjectionattack]) across key dimensions. Notably, Decepticon is the only benchmark that focuses on human-targeted adversarial elements, provides non-adversarial control pages for direct comparison, ensures archival determinism for reproducible offline evaluation, enforces avoidability constraints on dark patterns, includes a structured category taxonomy for adversarial patterns, employs dual metrics for success rate and dark pattern effectiveness, and systematically evaluates defense mechanisms. This comprehensive approach enables a more thorough understanding of dark pattern impacts on web agents compared to existing benchmarks and attack settings.
@@ Appendix I Comparison with Related Benchmarks and Attack Settings
Appendix J Modality Ablation
This section presents an ablation study on the input modalities, in order to quantify the contribution of different input modalities to dark pattern effectiveness. We conduct experiments using GLM-4.5V [vteam2025glm45vglm41vthinkingversatilemultimodal] on the Generated split of Decepticon, varying the input modalities provided to the model. We consider three configurations: (1) Accessibility Tree (a11y): only the accessibility tree representation of the webpage is provided as input; (2) Coordinates (coordinates): only the coordinates of clickable elements on the page are provided as input; (3) Screenshot + JS-based Set of Marks (SoM): both the screenshot of the webpage and the JS-based Set of Marks representation are provided as input. The results are summarized in Table 14.
@@
Appendix K Details on Human Study of Decepticon Tasks
@@
Appendix L Examples of Failures in Defenses
@@
L.1 In-Context Prompting Failure Examples
@@
L.2 Guardrail Failure Examples